### v1

@@ Abstract
The purpose of this paper is to propose a new multi-layer feedforward quaternion neural network model architecture, Reverse Quaternion Neural Network which utilizes the non-commutative nature of quaternion products, and to clarify its learning characteristics. While quaternion neural networks have been used in various fields, there has been no research report on the characteristics of multi-layer feedforward quaternion neural networks where weights are applied in the reverse direction. This paper investigates the learning characteristics of the Reverse Quaternion Neural Network from two perspectives: the learning speed and the generalization on rotation. As a result, it is found that the Reverse Quaternion Neural Network has a learning speed comparable to existing models and can obtain a different rotation representation from the existing models.
Keywords:
neural network learning quaternion non-commutative property rotation
1 Introduction
@@
In this paper, we propose a new feedforward quaternion neural network architecture, Reverse Quaternion Neural Networks (called RQNN for short here), which utilizes the non-commutative property of quaternions. In the model, by altering the order of quaternion multiplication, it is possible to construct neural networks with different characteristics. This approach is expected to capture information and offer a different expressive power that traditional quaternion neural networks cannot achieve. As a matter of facts , we conducted experiments on the learning speed and the generalization ability on rotation of the RQNN and the usual multi-layer feedforward quaternion neural network, and showed the differences between the two quaternion neural networks.
2 Reverse Quaternion Neural Network Architecture
@@ 2.1 Quaternion and Its Properties
From Eqs. (2.1) and (2.1), it is evident that the multiplication of quaternions is non-commutative. The conjugate quaternion of $q=a+b{\sf i}+c{\sf j}+d{\sf k}$, denoted as $\bar{q}$, is defined as
$\bar{q}=a-b{\sf i}-c{\sf j}-d{\sf k}.$ (6)
2.2 A Reverse Quaternion Neuron
@@
Figure 1: In the Reverse Quaternion Neural Network, unlike the usual H quaternion neural network, weights are applied in reverse order.
2.3 Reverse Quaternion Neural Network
@@
2.4 Learning Algorithm
Next, we derive a learning rule for the RQNN described in Section 2.3.
@@ 2.4 Learning Algorithm
$\Delta v_{nm}$ ::=:= $\Delta v_{nm}^{a}+\Delta v_{nm}^{b}{\sf i}+\Delta v_{nm}^{c}{\sf j}+\Delta v_{nm}^{d}{\sf k}$ (10)
$=$ $-\;\varepsilon\left(\frac{\partial E_{p}}{\partial v_{nm}^{a}}+\frac{\partial E_{p}}{\partial v_{nm}^{b}}{\sf i}+\frac{\partial E_{p}}{\partial v_{nm}^{c}}{\sf j}+\frac{\partial E_{p}}{\partial v_{nm}^{d}}{\sf k}\right),$
$\Delta\gamma_{n}$ ::=:= $\Delta\gamma_{n}^{a}+\Delta\gamma_{n}^{b}{\sf i}+\Delta\gamma_{n}^{c}{\sf j}+\Delta\gamma_{n}^{d}{\sf k}$ (11)
@@
$\Delta w_{ml}$ ::=:= $\Delta w_{ml}^{a}+\Delta w_{ml}^{b}{\sf i}+\Delta w_{ml}^{c}{\sf j}+\Delta w_{ml}^{d}{\sf k}$ (12)
$=$ $-\;\varepsilon\left(\frac{\partial E_{p}}{\partial w_{ml}^{a}}+\frac{\partial E_{p}}{\partial w_{ml}^{b}}{\sf i}+\frac{\partial E_{p}}{\partial w_{ml}^{c}}{\sf j}+\frac{\partial E_{p}}{\partial w_{ml}^{d}}{\sf k}\right),$
$\Delta\theta_{m}$ ::=:= $\Delta\theta_{m}^{a}+\Delta\theta_{m}^{b}{\sf i}+\Delta\theta_{m}^{c}{\sf j}+\Delta\theta_{m}^{d}{\sf k}$ (13)
$=$ $-\;\varepsilon\left(\frac{\partial E_{p}}{\partial\theta_{m}^{a}}+\frac{\partial E_{p}}{\partial\theta_{m}^{b}}{\sf i}+\frac{\partial E_{p}}{\partial\theta_{m}^{c}}{\sf j}+\frac{\partial E_{p}}{\partial\theta_{m}^{d}}{\sf k}\right)$
@@
$\Delta v_{nm}=$ $\ \bar{H}\Delta\gamma_{n},$ (14)
@@
$\Delta w_{ml}=$ $\ I_{l}{\Delta\bar{\theta}_{m}},$ (16)
$\Delta\theta_{m}=$ $\ (1-H^{a}_{m})H^{a}_{m}\cdot Re\left[\sum_{n}(\Delta{\bar{\gamma_{n}}v_{nm}})\right]$
$\quad+(1-H^{b}_{m})H^{b}_{m}\cdot Im^{i}\left[\sum_{n}(\Delta{\bar{\gamma_{n}}v_{nm}})\right]{\sf i}$
$\quad+(1-H^{c}_{m})H^{c}_{m}\cdot Im^{j}\left[\sum_{n}(\Delta{\bar{\gamma_{n}}v_{nm}})\right]{\sf j}$
$\quad+(1-H^{d}_{m})H^{d}_{m}\cdot Im^{k}\left[\sum_{n}(\Delta{\bar{\gamma_{n}}v_{nm}})\right]{\sf k}$ (17)
@@
$\Delta\theta_{m}$ $=$ $(1-H^{a}_{m})H^{a}_{m}\cdot Re\left[\sum_{n}(\Delta\gamma_{n}\overline{v}_{nm})\right]$ (21)
@@
The meaning of symbols in Eqs. (18) - (21) is the same as those in Eqs. (14) - (2.4). The differences between the learning rule of the RQNN and the one of the usual quaternion neural network are the positions of $\delta_{n}^{a},\delta_{n}^{b},\delta_{n}^{c},\delta_{n}^{d}$ in Eqs. (2.4) and (19), and the complex conjugate for $\Delta\gamma_{n},v_{nm}$ in Eqs. (2.4) and (21).
3 Experiments
@@
In order to investigate the learning characteristics of the RQNN, we conducted two experiments focusing on learning speed and rotation representation, and evaluated their performance.
@@
We call the multi-layer feedforward quaternion neural network where $xw$ is calculated for a weight $w\in\mathbb{H}$ and an input signal $x\in\mathbb{H}$ to a quaternion neuron QNN here for short. Also, we call the multi-layer feedforward real-valued neural network Real-NN here for short.
3.1 Learning Speed
First, we compared the learning speeds of the RQNN, the QNN, and the Real-NN. We used the learning data shown in Table 1, which was used in [7].
Table 1: Learning data[7]
Input Output
$1+{\sf i}+{\sf j}+{\sf k}$ $1$
$1+2{\sf i}+{\sf j}+2{\sf k}$ ${\sf i}$
$2+{\sf i}+2{\sf j}+{\sf k}$ ${\sf j}$
$2+2{\sf i}+2{\sf j}+2{\sf k}$ ${\sf k}$
To ensure fairness in computation, the Real-NN was structured as a 4-8-4 fully connected three-layer network, while both the QNN and the RQNN were structured as 1-6-1 fully connected three-layer networks (TABLE 2). Since one quaternion consists of four real numbers, it is calculated as four parameters.
Table 2: Layer configuration for each neural network and the total number of parameters.
Model Layer Total Params
Real–NN 4-8-4 76
QNN 1-6-1 76
RQNN 1-6-1 76
The sigmoid function was used for the real-valued function $f_{R}$ in the activation function (Eq. (9)). The initial first, second, third, and fourth parts of quaternion weights and quaternion biases were randomly generated from a uniform distribution ranging from $-0.3$ to $0.3$. The initial real-valued weights and real-valued biases of the Real-NN were also set the same way. The learning rate $\varepsilon$ was set to $0.1$. We determined that learning converged when the value of the learning error (loss) was below $0.01$. A total of 100 trials were performed. The average number of iterations required for convergence of 100 trials was evaluated. The results of the experiments are shown in Table 3. The typical learning curves are shown in Fig. 2.
It can be observed from Fig. 2 that despite having higher initial loss values, both the QNN and the RQNN converge more than twice as fast compared to the Real-NN. On the other hand, when comparing the QNN with the RQNN, their learning speeds are almost identical, and their learning curves follow the same trajectory.
Table 3: Average number of iterations required for convergence of 100 trials (Avg $\pm$ Std)
Real-NN QNN RQNN
 10701.98 $\pm$ 593.71 3111.84 $\pm$ 315.10 3306.22 $\pm$ 681.29
Figure 2: An example of learning curve. The yellow curve denotes the Real-NN, the blue curve the QNN, and the purple curve the RQNN. The bule and purple lines almost overlap.
3.2 Representation of Rotation
Next, we evaluated the generalization on rotation of the RQNN via a computer simulation: 1) rotation of lines, and 2) rotation of point cloud.
3.2.1 Rotation of lines
To investigate the generalization performance on rotation, 200 points were generated at equal intervals along each of the i, j, k axes in 3D space, ensuring that the angle between each data point was 90 degrees (Fig. 3). An experiment was then conducted to examine the behavior of the output of the RQNN for the test data after training. For comparison, the same experiment was also conducted with the traditional QNN. The input learning data were assigned to the points on the i-axis, the teacher data to the points on the j-axis, and the input test data to the points on the k-axis.
Figure 3: The blue represents the input learning data, the orange the teacher data, and the green the input test data.
Both the QNN and RQNN were configured as five-layer fully connected neural networks with layer sizes of 4-120-240-120-4. Since the data used in this experiment is linear (Fig. 3), the activation function was set to an identity function. The learning rate was set at 0.05, and training was performed until the training loss value fell below 0.01, with a total of 50 trials. The Adam optimization algorithm was utilized, and weights were initialized using the Xavier uniform distribution. Identical initial values were used for the weights of both the QNN and RQNN.
In order to quantitatively measure the differences in the outputs of the two networks after training, the angles formed by the outputs of both the QNN and the RQNN were calculated using Eq. (22). Additionally, the rotation angles from the input test data to the output data for each point in both the QNN and the RQNN were computed using the Euler angles (Eqs. (23) - (25)).
$\cos{\theta}$ $=\frac{\mathbf{v_{1}}\cdot\mathbf{v_{2}}}{\|\mathbf{v_{1}}\|\|\mathbf{v_{2}}\|}\ \ \mbox{where}\ \ \mathbf{v_{1}},\mathbf{v_{2}}\in\mathbb{R}^{3},$ (22)
roll $=\arctan 2(2(q_{w}q_{x}+q_{y}q_{z}),1-2(q_{x}^{2}+q_{y}^{2})),$ (23)
pitch =$=\begin{cases}\arcsin(2(q_{w}q_{y}-q_{z}q_{x})),&\!\!\!\!(\text{if }|2(q_{w}q_{y}-q_{z}q_{x})|\leq 1)\\ \frac{\pi}{2}\cdot\text{sign}(2(q_{w}q_{y}-q_{z}q_{x}))&(\text{otherwise}),\end{cases}$ (24)
yaw $=\arctan 2(2(q_{w}q_{z}+q_{x}q_{y}),1-2(q_{y}^{2}+q_{z}^{2}))$ (25)
for a quaternion $q=q_{w}+q_{x}{\sf i}+q_{y}{\sf j}+q_{z}{\sf k}\in\mathbb{H}$. Roll, pitch, and yaw, denoted as roll, pitch, and yaw, respectively, refer to rotations around the x-axis, y-axis, and z-axis. sign is a function that returns the sign of a number.
Fig. 4 displays a visualization of one of the experimental results. The blue line shows the output of the QNN for the input test data (red line) after training, and the purple line shows the output of the RQNN for the same input test data (red line) after training. We can find from Fig. 4 that the QNN and the RQNN output completely different values for the same input test data. Table 4 shows the average and the standard deviation of the angles between the outputs of the QNN and the RQNN, as well as those of the angles between the input test data and the outputs of the QNN over 50 trials, and those of the RQNN. We can find from Table 4 that the QNN and the RQNN learned to rotate about 90 degrees in different ways.
Table 5 shows the Euler angles (Eqs. (23) - (25)) of the rotation from the input test data to the outputs of the QNN and RQNN for the result shown in Fig. 4. It is also obvious from Table 5 that the QNN and the RQNN output completely different values for the same input test data.

### v2

@@ Abstract
The purpose of this paper is to propose a new neuron structure for multilayer feedforward quaternion neural networks called Reverse Quaternion Neural Network, which utilizes the non-commutative property of quaternion multiplication, and to clarify its learning characteristics. While quaternion neural networks have been used in various fields, there have been no research reports on the characteristics of multi-layer feedforward quaternion neural networks that apply weights in the reverse direction. In this paper, we examined the learning characteristics of the Reverse Quaternion Neural Network through two experimental studies: line rotation tasks and three-dimensional object rotation tasks using ModelNet10. Through comparative analysis with conventional models, we investigated both the generalization capabilities and internal learning dynamics.
Our results demonstrated that while the Reverse Quaternion Neural Network achieved performance comparable to existing models, it exhibited distinct characteristics in both rotational generalization capability and learning behavior, as evidenced by its gradient dynamics and feature acquisition process.
Index Terms:
neural network, learning, quaternion, non-commutative property, rotation
I Introduction
@@
In this paper, we proposed a new neuron structure for multi-layer feedforward quaternion neural networks called the Reverse Quaternion Neural Network (RQNN), which utilizes the non-commutative property of quaternions. The RQNN enables the construction of neural networks with different characteristics by altering the order of quaternion multiplication. This approach captures information and provides expressive capabilities that conventional quaternion neural networks cannot achieve. Specifically, we conducted experiments based on line rotation and three-dimensional object rotation tasks to evaluate the generalization capability and learning characteristics of the RQNN compared to conventional multilayer feedforward quaternion neural networks. Our experimental results revealed that the RQNN not only acquires rotational representations distinct from those of conventional quaternion neural networks but also exhibits unique learning characteristics.
II Reverse Quaternion Neural Network Architecture
@@ II-A Quaternion and Its Properties
From Eqs. (II-A) and (II-A), it is evident that the multiplication of quaternions is non-commutative. The conjugate quaternion of $q=a+b{\sf i}+c{\sf j}+d{\sf k}$, denoted as $\overline{q}$, is defined as
$\overline{q}=a-b{\sf i}-c{\sf j}-d{\sf k}.$ (6)
II-B A Reverse Quaternion Neuron
@@
Figure 1: In the Reverse Quaternion Neural Network, unlike the usual quaternion neural network, weights are applied in reverse order.
II-C Reverse Quaternion Neural Network
@@
II-D Learning Algorithm
Next, we derive a learning rule for the RQNN described in Section II-C.
@@ II-D Learning Algorithm
$\Delta v_{nm}$ $:=$ $\Delta v_{nm}^{a}+\Delta v_{nm}^{b}{\sf i}+\Delta v_{nm}^{c}{\sf j}+\Delta v_{nm}^{d}{\sf k}$ (10)
$=$ $-\;\varepsilon\left(\frac{\partial E_{p}}{\partial v_{nm}^{a}}+\frac{\partial E_{p}}{\partial v_{nm}^{b}}{\sf i}+\frac{\partial E_{p}}{\partial v_{nm}^{c}}{\sf j}+\frac{\partial E_{p}}{\partial v_{nm}^{d}}{\sf k}\right),\ \$
$\Delta\gamma_{n}$ $:=$ $\Delta\gamma_{n}^{a}+\Delta\gamma_{n}^{b}{\sf i}+\Delta\gamma_{n}^{c}{\sf j}+\Delta\gamma_{n}^{d}{\sf k}$ (11)
@@
$\Delta w_{ml}$ $:=$ $\Delta w_{ml}^{a}+\Delta w_{ml}^{b}{\sf i}+\Delta w_{ml}^{c}{\sf j}+\Delta w_{ml}^{d}{\sf k}$ (12)
$=$ $-\;\varepsilon\left(\frac{\partial E_{p}}{\partial w_{ml}^{a}}+\frac{\partial E_{p}}{\partial w_{ml}^{b}}{\sf i}+\frac{\partial E_{p}}{\partial w_{ml}^{c}}{\sf j}+\frac{\partial E_{p}}{\partial w_{ml}^{d}}{\sf k}\right),\$
$\Delta\theta_{m}$ $:=$ $\Delta\theta_{m}^{a}+\Delta\theta_{m}^{b}{\sf i}+\Delta\theta_{m}^{c}{\sf j}+\Delta\theta_{m}^{d}{\sf k}$ (13)
$=$ $-\;\varepsilon\left(\frac{\partial E_{p}}{\partial\theta_{m}^{a}}+\frac{\partial E_{p}}{\partial\theta_{m}^{b}}{\sf i}+\frac{\partial E_{p}}{\partial\theta_{m}^{c}}{\sf j}+\frac{\partial E_{p}}{\partial\theta_{m}^{d}}{\sf k}\right).$
@@
$\Delta v_{nm}=$ $\ \overline{H}_{m}\Delta\gamma_{n},$ (14)
@@
$\Delta w_{ml}=$ $\ I_{l}{\Delta\overline{\theta}_{m}},$ (16)
$\Delta\theta_{m}=$ $\ (1-H^{a}_{m})H^{a}_{m}\cdot Re\left[\sum_{n}(\Delta{\overline{\gamma_{n}}v_{nm}})\right]$
$\quad+(1-H^{b}_{m})H^{b}_{m}\cdot Im^{i}\left[\sum_{n}(\Delta{\overline{\gamma_{n}}v_{nm}})\right]{\sf i}$
$\quad+(1-H^{c}_{m})H^{c}_{m}\cdot Im^{j}\left[\sum_{n}(\Delta{\overline{\gamma_{n}}v_{nm}})\right]{\sf j}$
$\quad+(1-H^{d}_{m})H^{d}_{m}\cdot Im^{k}\left[\sum_{n}(\Delta{\overline{\gamma_{n}}v_{nm}})\right]{\sf k}.$ (17)
@@
$\Delta\theta_{m}$ $=$ $(1-H^{a}_{m})H^{a}_{m}\cdot Re\left[\sum_{n}(\Delta\gamma_{n}\overline{v}_{nm})\right]$
@@
The meaning of symbols in Eqs. (18) - (II-D) is the same as those in Eqs. (14) - (II-D). The differences between the learning rule of the RQNN and the one of the usual quaternion neural network are the positions of $\delta_{n}^{a},\delta_{n}^{b},\delta_{n}^{c},\delta_{n}^{d}$ in Eqs. (II-D) and (19), $\Delta\theta_{m},I_{l}$ in Eqs. (16) and (20), and the quaternion conjugate for $\Delta\gamma_{n},v_{nm}$ in Eqs. (II-D) and (II-D).
III Experiments
@@
In order to investigate the learning characteristics of the RQNN, we conducted experiments focusing on two rotation representations and evaluated their performance.
@@
We call the multi-layer feedforward network with quaternion neurons, where $xw$ is calculated for a weight $w\in\mathbb{H}$ and an input signal $x\in\mathbb{H}$, QNN for short.
III-A Rotation of Lines Task
We evaluated the RQNN’s generalization performance for rotations through line rotation simulations. In the three-dimensional space defined by the i, j, and k axes, we set a reference line along the diagonal direction $1,1,1]/\sqrt{3}$ and sampled 99 points at 0.01 intervals from 0.01 to 0.99. This was used as the input test data. Furthermore, we created a line along $1,-2,1]/\sqrt{6}$ that is perpendicular to the reference line, which served as the output target data, and another line along $1,0,-1]/\sqrt{2}$ as the input learning data (Fig. 2).
Figure 2: The blue represents the input learning data, the yellow the output target data, and the red the input test data. The figure on the right is a view of the figure on the left from a different angle.
Both the RQNN and the QNN models were implemented as fully connected three-layer neural networks with a 4-8-4 architecture. The weights of the QNN and the RQNN were initialized from the same standard normal distribution to ensure a fair comparison. The sigmoid function was employed as the activation function for the hidden layer, while the identity function was used for the output layer. The learning rate was set to $5\times 10^{-3}$, and training was continued until the training loss value decreased below 0.05. To assess the learning characteristics of both models, we evaluated the differences in their outputs using the following four distinct evaluation metrics. The primary objective of this research is not to compare performance on test data, but rather to analyze the differences in internal learning characteristics resulting from different quaternion multiplication orders. Therefore, instead of conventional accuracy metrics, we employed these metrics to capture the differences in the internal representations of the models.
  1. 1.
    Angle Differences: The angle between the tangent vectors at each point of the output curves of the QNN and the RQNN is calculated and converted from radians to degrees through multiplication by $180^{\circ}/\pi$, as defined by Eqs.(22) and (23).
    $\theta_{i}=\arccos\!\Bigl{(}\frac{\mathbf{t}^{Q}_{i}\cdot\mathbf{t}^{RQ}_{i}}{\|\mathbf{t}^{Q}_{i}\|\|\mathbf{t}^{RQ}_{i}\|}\Bigr{)},$ (22)
    $\text{AngleDiff}=\frac{180^{\circ}}{\pi(N-1)}\sum_{i=1}^{N-1}\theta_{i}.$ (23)
    Here, $\mathbf{Q}_{i},\mathbf{RQ}_{i}\in\mathbb{R}^{3}$ denote the position vectors of the $i$-th output point of the QNN and the RQNN respectively, where $i\in\{1,\ldots,N\}$ and $N$ is the total number of points. The tangent vectors $\mathbf{t}^{Q}_{i},\mathbf{t}^{RQ}_{i}\in\mathbb{R}^{3}$ are defined at the $i$-th point as $\mathbf{t}^{Q}_{i}=\mathbf{Q}_{i+1}-\mathbf{Q}_{i}$ for the QNN and $\mathbf{t}^{RQ}_{i}=\mathbf{RQ}_{i+1}-\mathbf{RQ}_{i}$ for the RQNN.
  2. 2.
    Trajectory Differences: The distance between the two curves is calculated using the Euclidean distance, as defined by Eq.(2), quantifying the overall difference in the curve shapes.
    $\text{TrajDiff}=\frac{1}{N}\sum_{i=1}^{N}\|\mathbf{Q}_{i}-\mathbf{RQ}_{i}\|$
    $\qquad+0.1\cdot\frac{1}{N-2}\sum_{i=1}^{N-2}|\kappa^{Q}_{i}-\kappa^{RQ}_{i}|.$ (24)
    $\text{where}\>\kappa^{m}_{i}=\frac{\|\mathbf{v}^{m}_{i}\times\mathbf{a}^{m}_{i}\|}{\|\mathbf{v}^{m}_{i}\|^{3}},\quad m\in\{Q,RQ\}$
    Here, $\kappa^{m}_{i}\in\mathbb{R}$ denotes the curvature at the $i$-th point for model $m\in\{Q,RQ\}$. The first derivative $\mathbf{v}^{m}_{i}\in\mathbb{R}^{3}$ is defined as $\mathbf{v}^{m}_{i}=\mathbf{Q}_{i+1}-\mathbf{Q}_{i}$ for $m=Q$ and $\mathbf{v}^{m}_{i}=\mathbf{RQ}_{i+1}-\mathbf{RQ}_{i}$ for $m=RQ$. The second derivative $\mathbf{a}^{m}_{i}\in\mathbb{R}^{3}$ is computed as $\mathbf{a}^{m}_{i}=\mathbf{v}^{m}_{i+1}-\mathbf{v}^{m}_{i}$.
  3. 3.
    Phase Space Differences: The differences in the first and second derivatives between the QNN and the RQNN are evaluated as defined by Eq.(3). This metric combines the differences in first derivatives with a weighted term for second derivatives to quantify the dynamic characteristics of the curves.
    PhaseDiff $=\frac{1}{N-1}\sum_{i=1}^{N-1}\|\mathbf{v}^{Q}_{i}-\mathbf{v}^{RQ}_{i}\|$
    $\qquad+0.1\cdot\frac{1}{N-2}\sum_{i=1}^{N-2}\|\mathbf{a}^{Q}_{i}-\mathbf{a}^{RQ}_{i}\|$ (25)
  4. 4.
    Geometric Consistency: The deviation of the angles between the tangent vectors from 90 degrees is calculated as defined by Eq.(4). The angle calculation in radians is converted to degrees through multiplication by $180^{\circ}/\pi$, measuring the geometric relationship between the test data and the outputs of both models.
    GeoConsist $=\frac{1}{2}\Bigg{\{}\sum_{m\in\{Q,RQ\}}\frac{1}{N-1}\sum_{i=1}^{N-1}$
    $\quad\left|\arccos\!\Bigl{(}\frac{\mathbf{v}^{m}_{i}\cdot\mathbf{t}_{test,i}}{\|\mathbf{v}^{m}_{i}\|\,\|\mathbf{t}_{test,i}\|}\Bigr{)}\cdot\frac{180^{\circ}}{\pi}-90^{\circ}\right|\Bigg{\}}$ (26)
    Here, $\mathbf{t}_{test,i}$ denotes the tangent vector of the test data at the $i$-th point, calculated as $\mathbf{t}_{test,i}=\mathbf{T}_{i+1}-\mathbf{T}_{i}$, where $\mathbf{T}_{i}\in\mathbb{R}^{3}$ represents the position vector of the $i$-th point on the test data curve.
Fig. 3 shows a representative example of the experimental results. The blue and purple lines represent the outputs of the QNN and the RQNN respectively when the test data (shown in red) is provided as input after training. As shown in Fig. 3, the QNN and the RQNN generate completely different output patterns despite being given the same test data as input. Fig. 4 shows the evolution of four evaluation metrics over 50 trials for both the QNN and the RQNN outputs, computed using Eqs.(22) - (4). Table I summarizes the statistical analysis results of these four evaluation metrics across the 50 trials.
As shown in Figs. 3 and 4 and Table I, the QNN and the RQNN exhibited clearly different learning behaviors, despite having identical initial weights and test data. Notably, the angle differences between the models’ tangent vectors demonstrated a significant discrepancy, averaging 87.65 degrees.